Company A makes coffee machines which are supplied to restaurants in boxes. The weekly demand for the coffee machines from the distributors is 1000 boxes. Company A works two (2) 8-hour shifts daily, 5 days a week. In each shift the workers are given a 30-minute paid lunch break during which no production occurs. The takt time for scalpel boxes is:
Answer #1

Takt time = total working time/ total demand

Given data

Weekly demand for coffee machines = 1000 boxes

SO, total demand = 1000 boxes

Working hours per day = 8 hours

Working hours per week = 8*5 = 40 hours

Working hours per week in minutes = 40*60 = 2400 minutes

Break hours per week for lunch = 5*30 =150

Total productive working hours per week = 2400-150 = 2250 minutes

Takt time = Total productive working hours per week/ total weekly demand

= 2250/ 1000

= 2.25 minutes